Bramble Decorating uses a job order cost system to collect the costs of its interior decorating business. Each client’s consultation is treated as a separate job. Overhead is applied to each job based on the number of decorator hours incurred. Listed below are data for the current year.

Estimated overhead $1,000,000
Actual overhead $994,800
Estimated decorator hours 40,000
Actual decorator hours 38,000


The company uses the account Operating Overhead in place of Manufacturing Overhead and the account Service Contracts in Process in place of Work in Process.

Compute the predetermined overhead rate.

Predetermined overhead rate

$

per decorator hour

Prepare the entry to apply the overhead for the year.

Determine whether the overhead was under- or overapplied and by how much.

Overhead $                                                                       Overapplied or Underapplied?
Answer #1

a) Predetermine overhead rate = 1000000/40000 = $25 per hour

b) Applied overhead = 38000*25 = 950000

Actual overhead = 994800

Overhead (994800-950000) = 44800 Under applied